two integers with a difference of -11and a sum of -3 what are the two numbers?
Answer by ewatrrr(24785)

Hi,
*Hint: Let x and y represent the two integers
The question states the following to be true:
.
x + y = -3
x - y = -11
.
Solve the system using addition.
2*x = -14
x = -7
.
(y = -x-3)
y = 4
.
Two numbers are -7 and 4
.
Check Your answer:
-7 + 4 = -3